Problem lampoly SNF 2009-04-17 16:18:56: HBr

mtang at mtang at
Fri Apr 17 16:18:57 PDT 2009

The HBr problem from yesterday we suspect might be cross-sensitivity of the TGO sensor with some environmental conditions.  It was quiet overnight and much less noisy today.  HBr has been turned back on.  But please watch your HBr flows when you process critical material.  (If the sensor goes off in the vault, there will be no audible alarm in the lab, but the HBr will shut off.)

More information about the lampoly-pcs mailing list